Trains A and B are going in the same direction on parallel tracks. Train A is moving at 40 miles per hour. Train B is moving at 60 miles per hour. Train A passes a station at 2:15am. If Train B passes the same station at 2:30am, what time will Train B catch up to Train A ?
Answer by sudhanshu_kmr(1152)


speed of A is 40 and speed of B is 60 miles/h.
relative velocity = 60 -40 = 20 miles/h.

As train A passes the station before 15 minutes than B, so we firstly find the
distance traveled by A during this time = 40/4 = 10 miles (15 minutes = 1/4 hour)

now we find time when B will catch A = distance/speed
= 10/20 = 1/2 hour
1/2 hour after 2:30 is 3:00

so, B will catch the A at 3:00 am.

Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......